5. Given ΔG =-16.5 kJ/mol for the reaction ½ N2 + 3/2 H2 → NH3, find the equilibrium constant for: (a) the above reaction as written (b) the reaction N2+ 3H2 → 2NH3
Answer #1

a)

T = 298.15 K

deltaG = -16.5 KJ/mol

deltaG = -16500 J/mol

we have below equation to be used:

deltaG = -R*T*ln Kc

-16500 = - 8.314*298.15* ln(Kc)

ln Kc = 6.6564

Kc = 778

Answer: 778

b)

This reaction is 2 times of above reaction

So,

Kc2 = Kc1^2

= (778)^2

= 6.05*10^5

Answer: 6.05*10^5